# Large Continual Instruction Assistant

### v3

@@ 5.2 Continual Instruction Tuning Results
Comparison to SOTA: Based on LLaVA-7B, we compare the performance in Table 1. We observe that our method can improve the best of other methods (EProj method) by +3.85@Avg.ACC, -3.49@Forgetting, and +0.79@New.ACC, demonstrating its excellent anti-forgetting and continual tuning ability. Although methods like LWF (Li & Hoiem, 2017) and EWC (Kirkpatrick et al., 2017) can resist forgetting, their plasticity is greatly influenced. PGP (Qiao et al., 2024a), MT (Zhu et al., 2024) and EProj (He et al., 2023) can perform well both in stability and plasticity, while the Avg.ACC still needs to be improved. It is highlighted that our method owns the highest Avg.ACC, the lowest forgetting and the comparative higher New.ACC among these methods, which shows that our method can achieve the best trade-off between plasticity and stability.
Furthermore, based on LLaVA-13B, we compare the continual instruction tuning performance in Table 2 on larger MLLM. We observe that our method can improve the Avg.ACC (+3.68), the New.ACC (+0.59) and reduce the Forgetting (-3.53) compared with the best of other methods (EProj method), highlighting its superior anti-forgetting and continual tuning capabilities. When compared with other SOTA methods, e.g. PGP (Qiao et al., 2024a), MT (Zhu et al., 2024) and EProj (He et al., 2023), it is also noteworthy that our method owns the highest Avg.ACC, the lowest forgetting and the comparative higher New.ACC among these methods, which shows that our method can achieve the best trade-off between plasticity and stability in larger MLLM and lays the foundation for the application of our method to MLLM with more parameters and stronger capabilities.
@@ 5.3 Robust Performance
Table 6: Ablation study results.
@@ 5.4 Analysis of Examples
In Figure 7, we show three testing instances with outputs from distinct methods and observe that our method can revise errors caused by continual instruction tuning. For the GQA instance (as shown in the left part), our method can keep the geographic location recognition capability of MLLMs after tuning. LoRA Fine-Tune seems to lose such ability and mistakes the chair for table. Although MT and EProj can recognize the object, their outputs are still not aligned with the ground truth, i.e. the first letter is not capitalized and the plural form is used. For the ImageNet instance (as shown in the middle part), our method can reserve the fine-grained knowledge of MLLMs after tuning. In contrast, LoRA Fine-Tune can only recognize the coarse attribute of the object. MT even appears the hallucination. For the Grounding instance (as shown in the right part), our method can protect the visual grounding ability in real scenarios. It is obvious that the IoU between our output and ground truth is the largest compared to other methods. Additionally, we are surprised to find that our method can spontaneously suppress the occurrence of hallucinations appearing in the continual instruction tuning. (Zhai et al., 2023) deem that the hallucination in large models is related to the catastrophic forgetting in continual tuning. The above view is consistent with our multiple rounds of dialogue results, which are shown in Appendix D.
@@ Appendix A Decompose of EMA Update
Observing the equation form, based on the method of summarization and induction, we have the following assumption for iteration $n-1$:
@@ Appendix C $\beta_{t}$ Solving Process
$10=\mathcal{L}^{\prime}(\theta_{t})+\frac{\beta_{t}}{(\beta_{t}-1)}\mathcal{L}^{\prime\prime}(\theta_{t})\Delta\theta+$ 1 . (36)
@@
Substitute 37 into Eq.(36):
$10=\mathcal{L}^{\prime}(\theta_{t})+\beta_{t}(\theta_{t-1}^{*}-\theta_{t})\mathcal{L}^{\prime\prime}(\theta_{t})+$ 1 . (38)
@@ Appendix D Cases of Multiple Rounds of Dialogue
In this section, we test the zero-shot performance of MLLMs continually fine-tuned with our method on the multiple rounds of dialogue tasks. Images and questions are from (Liu et al., 2024b). To have a comparison, we also test the zero-shot performance of MLLMs continually fine-tuned with the baseline on the multiple rounds of dialogue tasks.
@@ Appendix F Influence of EMA summation weight
We conduct a toy experiment to discuss that distinct stable EMA weights can greatly influence the continual instruction tuning results. We choose six Visual Question Answering (VQA) tasks, including ScienceQA (Lu et al., 2022), TextVQA(Singh et al., 2019), GQA (Hudson & Manning, 2019), VizWiz (Gurari et al., 2018), (Goyal et al., 2017) and (Mishra et al., 2019). Based on the LLaVA-7B backbone, we fine-tune on the six tasks with the above orders. To have a comparison, we utilize three EMA weights and the results are shown in Table 7
Table 7: Continual Instruction Tuning Performance of Distinct EMA Weights
@@
Appendix G Comparions Between Our Method With Others
Notice that existing methods e.g. L2P (Wang et al., 2022b) and EProj (He et al., 2023) also own the training parameters (Prompt or Linear Projection Layer) selection mechanism, which are similar to our instruction grouping and parameters expansion strategy. Thus, in this section, we analyze and illustrate the differences between our method and theirs.
L2P owns a fixable prompt pool, which consists of key-prompt pairs. During the training and testing stage, L2P adopts a frozen pre-trained Vision-Transformer to extract the image embedding and match it with the keys by using cosine similarity. After matching, it would find the prompts with the top-K similarity. Notice here, besides the training cost of prompt parameters, key parameters also need to be optimized in the training iteration.
EProj proposes a novel task-similarity model expansion, which decides whether to retrain or expand the image-to-text projection layer by measuring the cosine similarity between current task and history tasks. To be specific, it would collect the mean embeddings of image $e(v)$, instruction $e(t)$, and output $e(o)$ of the entire dataset to comprise the task embeddings. The corresponding embeddings are encoded by the BERT model and the frozen Vision-Transformer in MLLMs respectively. During the training stage, it also extra trains key parameters to match with the input task embeddings aiming to retrieve the corresponding projection layer at the testing time.
@@ Appendix H Term Frequency-Inverse Document Frequency
$TF(t,d)=\frac{\text{Number of occurrences of term }t\text{ in document }d}{\text{Total number of terms in document }d}$ (46)
Inverse Document Frequency reduces the weight of terms that occur in many documents, emphasizing terms unique to fewer documents:
$IDF(t,D)=\log\frac{|D|}{1+|\{d\in D:t\in d\}|}$ (47)
where:
  • $|D|$: Total number of documents in the corpus.
  • |$|\{d\in D:t\in d\}|$: Number of documents containing term $t$.
@@
$TF\text{-}IDF(t,d,D)=TF(t,d)\cdot IDF(t,D)$ (48)
TF-IDF is effective for capturing term importance and extracting feature for text classification, making it a simple yet powerful tool in text processing. In our code, we utilize the TfidfVectorizer class in sklearn library to tokenize the instruction texts into numerical vector.
@@ Appendix I Instruction Grouping Results
In this section, we summarize the details of three kinds of instruction templates, as shown in Table O. The instruction grouping results are shown in Table 10. It can be observed that in the three kinds of instruction templates, our strategy divides the instructions of the eight tasks into 4 groups (both for Origin and 10Type), 5 groups (for Diverse), respectively, which can be deemed as a limited expansion compared to the total number of tasks.
@@
In the implementation of our method, the codebase is based on CoIN (Chen et al., 2024a). The inserted LoRA in each module layer of LLM has a rank of 128. For each fine-tuning dataset, the training epoch is set to 1, and the initial learning rate and weight decay are configured at 2e-4 and 0. The max length of input text is fitted as 2048. Additionally, we adopt gradient checkpoint strategy and mixed precision mode of TF32 and BF16. Furthermore, we also utilize the ZeRO stage: 0 mode of DeepSpeed for training. All experiments are conducted on 8 NVIDIA H100 GPUs with 80GB of memory.
@@
LoRA Fine-Tune (Hu et al., 2022) prepends LoRA parameter efficient tuning paradigm into LLM. In the training stage, it only trains the linear projector and LoRA parameters, with frozen vision encoder and LLM; MoELoRA (Chen et al., 2024a) is based on the LoRA, and the number of experts for each MoE layer is set to 2; LWF (Li & Hoiem, 2017) calculates the results of the new dataset samples on both the old and new models. After that, it calculates the distillation loss and adds it to the loss function as a regularization penalty term. EWC (Kirkpatrick et al., 2017) considers the change of the training parameters and proposes the specific parameters changing loss as a regularization penalty. PGP (Qiao et al., 2024a) introduces a gradient projection method for efficient parameters, and changes the gradient direction orthogonal to the previous feature subspace. E-Proj (He et al., 2023) is the representative SOTA dynamic modal method, which expands the visual projection layer with the task increasing. MT (Zhu et al., 2024) is the representative SOTA regularization method, which compensates for changes in trainable parameters.
@@
It is worth noting that our judgment of whether the prediction results are correct or not is strictly based on the direct comparison between outputs of MLLMs and ground truths, which is defined as Truth Alignment in (Chen et al., 2024a). Therefore, our judgment criteria would be more stringent.
@@
New Accuracy (New.ACC) is employed to average the test accuracy of new datasets, which refers to the plasticity performance.
@@
Average $\text{Average Accuracy}=\frac{1}{T}\sum_{i=1}^{T}A_{T,i},$ (49)
$\text{Forgetting}=\frac{1}{T-1}\sum_{i=1}^{T-1}{A_{T,i}–\text{max}(A_{j,i})_{j\in[i,T-1]}},$ (50)
New $\text{New Accuracy}=\frac{1}{T}\sum_{i=1}^{T}A_{i,i},$ (51)
@@
FT-init 35.70.2 18.50.7 --4.60.2
L2 35.60.1 17.50.5 --3.81.2
EWC 34.50.6 16.80.4 --6.81.5
AGEM (10) 33.20.4 19.10.1 --7.31.0
AGEM (50) 34.90.9 18.11.0 --6.00.9
Replay (10) 38.40.7 23.70.0 --1.30.5
@@
Dataset. We use the InstrDialog Stream dataset from (Zhang et al., 2023), which consists of 4 tasks from dialogue state tracking, 11 tasks from dialogue generation, and 4 tasks from intent identification, resulting in a total of 19 dialogue tasks.
Model. We use the instruction-tuned T5 model from (Zhang et al., 2023), which has learned to understand some instructions and can act as a good starting point to conduct subsequent learning.
Comparison Methods. (1). L2 and EWC regularization methods, both of which leverage a Fisher information matrix to mitigate forgetting by regularizing the loss function, thereby penalizing changes to the crucial parameters of previous tasks. (2). Replay method, which stores random instances from each task in a memory buffer, then trains the model jointly on both the new task data and the stored instances. (3). AGEM method, which adds constraints to prevent parameter update from increasing the loss of each previous task. The loss for these tasks is computed based on the instances saved in the memory. (4). We adopt AdapterCL, which freezes the pre-trained model and trains an independent Adapter for each task. (5). For ours, we freeze the pre-trained model and train a set of Adapters. Specifically, we utilize the instruction grouping strategy to decide whether to retrain or expand the Adapter according to the current task instructions and adopt dynamical EMA update when retraining the Adapter. Notice here, besides AdapterCL and ours, which freeze the LLM and only train the Adapter, others adopt the full fine-tuning (FFT) mode.
Metrics. Following (Zhang et al., 2023), we use the following metrics to measure the continual tuning performance. Let $a_{j,i}$ be the ROUGE-L score of the model on the test set of task $t_{i}$ after training on task $t_{j}$, we define the following:
@@
$\textbf{\text{AR}}=\frac{1}{T}\sum_{i=1}^{T}a_{T,i}$ (52)
@@
$\textbf{\text{FWT}}=\frac{1}{T-1}\sum_{i=2}^{T}a_{i-1,i}$ (53)
@@
$\textbf{\text{BWT}}=\frac{1}{T-1}\sum_{i=1}^{T-1}\left(a_{T,i}-a_{i,i}\right)$ (54)
@@
In order to validate the robustness of our method, we adopt the following three types of training instructions. For detailed instruction templates please kindly refer to Table O.
1). Original instruction type: Each task owns only one instruction, and several tasks share the same instruction.
2). Diverse instruction type: Each task owns only one instruction, and different tasks are specifically tailored to distinct instructions.
@@
ScienceQA Answer with the option’s letter from the given choices directly Answer with the option’s letter from the given choices directly Answer with the option’s letter from the given choices directly Select the correct answer from the given choices and respond with the letter of the chosen option Determine the correct option from the provided choices and reply with its corresponding letter Pick the correct answer from the listed options and provide the letter of the selected option Identify the correct choice from the options below and respond with the letter of the correct option From the given choices, choose the correct answer and respond with the letter of that choice Choose the right answer from the options and respond with its letter Select the correct answer from the provided options and reply with the letter associated with it From the given choices, select the correct answer and reply with the letter of the chosen option Identify the correct option from the choices provided and respond with the letter of the correct option From the given choices, pick the correct answer and respond by indicating the letter of the correct option
[…]

### v5

@@ 5.2 Continual Instruction Tuning Results
Comparison to SOTA: Based on LLaVA-7B, we compare the performance in Table 1. We observe that our method can improve the best of other methods (EProj method) by +3.85@Avg.ACC, -3.49@Forgetting, and +0.79@New.ACC, demonstrating its excellent anti-forgetting and continual tuning ability. Although methods like LWF (lwf) and EWC (ewc) can resist forgetting, their plasticity is greatly influenced. PGP (pgp), MT (mt) and EProj (mcit) can perform well both in stability and plasticity, while the Avg.ACC still needs to be improved. It is highlighted that our method owns the highest Avg.ACC, the lowest forgetting and the comparative higher New.ACC among these methods, which shows that our method can achieve the best trade-off between plasticity and stability.
Furthermore, based on LLaVA-13B, we compare the continual instruction tuning performance in Table 2 on larger MLLM. We observe that our method can improve the Avg.ACC (+3.68), the New.ACC (+0.59) and reduce the Forgetting (-3.53) compared with the best of other methods (EProj method), highlighting its superior anti-forgetting and continual tuning capabilities. When compared with other SOTA methods, e.g. PGP (pgp), MT (mt) and EProj (mcit), it is also noteworthy that our method owns the highest Avg.ACC, the lowest Forgetting and the comparative higher New.ACC among these methods, which shows that our method can achieve the best trade-off between plasticity and stability in larger MLLM and lays the foundation for the application of our method to MLLM with more parameters and stronger capabilities.
@@ 5.3 Robust Performance
Table 6: Ablation study results for each proposed component.
@@ 5.4 Analysis of Examples
In Figure 7, we show three testing instances with outputs from distinct methods and observe that our method can revise errors caused by continual instruction tuning. For the GQA instance (as shown in the left part), our method can keep the geographic location recognition capability of MLLMs after tuning. LoRA Fine-Tune seems to lose such ability and mistakes the chair for table. Although MT and EProj can recognize the object, their outputs are still not aligned with the ground truth, i.e. the first letter is not capitalized and the plural form is used. For the ImageNet instance (as shown in the middle part), our method can reserve the fine-grained knowledge of MLLMs after tuning. In contrast, LoRA Fine-Tune can only recognize the coarse attribute of the object. MT even appears the hallucination. For the Grounding instance (as shown in the right part), our method can protect the visual grounding ability in real scenarios. It is obvious that the IoU between our output and ground truth is the largest compared to other methods. Additionally, we are surprised to find that our method can spontaneously suppress the occurrence of hallucinations in the continual instruction tuning. (hal) deem that the hallucination in large models is related to the catastrophic forgetting in continual tuning. The above view is consistent with our multiple rounds of dialogue results, which are shown in Appendix D.
@@ 6 Conclusion
7 Impact Statement
1.The potential of our method to enhance LFMs’ continual instruction tuning performance and mitigate catastrophic forgetting. 2.Its relevance to real-world applications, such as lifelong AI assistants and LLM continual evolution. 3.Ethical considerations, including the reduction of privacy risks through the adoption of a non-data-rehearsal approach.
@@ Appendix A Decompose of EMA Update
Observing the equation form, based on the method of summarization and induction, we have the following assumption for iteration n1n-1:
@@ Appendix C $\beta_{t}$ Solving Process
$0=\mathcal{L}^{\prime}(\theta_{t})+\frac{\beta_{t}}{(\beta_{t}-1)}\mathcal{L}^{\prime\prime}(\theta_{t})\Delta\theta+1.$ (36)
@@
Substitute Eq.(37) into Eq.(36):
$0=\mathcal{L}^{\prime}(\theta_{t})+\beta_{t}(\theta_{t-1}^{*}-\theta_{t})\mathcal{L}^{\prime\prime}(\theta_{t})+1.$ (38)
@@ Appendix D Cases of Multiple Rounds of Dialogue
In this section, we test the zero-shot performance of MLLMs continually fine-tuned with our method on the multiple rounds of dialogue tasks. Images and questions are from (llava). To have a comparison, we also test the zero-shot performance of MLLMs continually fine-tuned with the baseline on the multiple rounds of dialogue tasks.
@@ Appendix F Influence of EMA summation weight
We conduct a toy experiment to discuss that distinct stable EMA weights can greatly influence the continual instruction tuning results. We fine-tune on six tasks, including ScienceQA, TextVQA, GQA, VizWiz, VQAv2, and OCRVQA, based on the LLaVA-7B backbone. To have a comparison, we utilize three EMA weights and the results are shown in Table 7.
Table 7: Continual Instruction Tuning Performance of Distinct EMA Weights.
@@
Appendix G Comparisons Between Our Method With Others
Notice that existing methods e.g. L2P (l2p) and EProj (mcit) also own the training parameters (Prompt or Linear Projection Layer) selection mechanism, which are similar to our instruction grouping and parameters expansion strategy. Thus, in this section, we analyze and illustrate the differences between our method and theirs.
L2P owns a fixed prompt pool, which consists of key-prompt pairs. During the training and testing stage, L2P adopts a frozen pre-trained Vision-Transformer to extract the image embedding and match it with the keys by using cosine similarity. After matching, it would find the prompts with the top-K similarity. Notice here, besides the training cost of prompt parameters, key parameters also need to be optimized in the training iteration.
EProj proposes a novel task-similarity model expansion, which decides whether to retrain or expand the image-to-text projection layer by measuring the cosine similarity between the current task and history tasks. To be specific, it would collect the mean embeddings of image $e(v)$, instruction $e(t)$, and output $e(o)$ of the entire dataset to comprise the task embeddings. The corresponding embeddings are encoded by the BERT model and the frozen Vision-Transformer in MLLMs, respectively. During the training stage, it also trains key parameters to match with the input task embeddings, aiming to retrieve the corresponding projection layer at the testing time.
@@ Appendix H Term Frequency-Inverse Document Frequency
[…]
Inverse Document Frequency reduces the weight of terms that in many documents, emphasizing terms in fewer documents:
$IDF(t,D)=\log\frac{|D|}{1+|\{d\in D:t\in d\}|},$ (47)
where $|D|$ is total number of documents in the corpus and $|\{d\in D:t\in d\}|$ is number of documents containing term $t$.
@@
$TF\text{-}IDF(t,d,D)=TF(t,d)\cdot IDF(t,D).$ (48)
TF-IDF is effective for capturing term importance and extracting features for text classification, making it a simple yet powerful tool in text processing. In our code, we utilize the TfidfVectorizer class in sklearn library to tokenize the instruction texts into numerical vector.
@@ Appendix I Instruction Grouping Results
In this section, we summarize the details of three kinds of instruction templates, as shown in Table 9. The instruction grouping results are shown in Table 10. It can be observed that in the three kinds of instruction templates, our strategy divides the instructions of the eight tasks into 4 groups (both for Origin and 10Type), 5 groups (for Diverse), respectively, which can be deemed as a limited expansion compared to the total number of tasks.
@@
In the implementation of our method, the codebase is based on CoIN (coin). The inserted LoRA in each module layer of LLM has a rank of 128. For each fine-tuning dataset, the training epoch is set to 1, and the initial learning rate and weight decay are configured at 2e-4 and 0. The max length of input text is fitted as 2048. Additionally, we adopt gradient checkpoint strategy and mixed precision mode of TF32 and BF16. Furthermore, we also utilize the ZeRO stage: 0 mode of DeepSpeed for training.
@@
LoRA Fine-Tune (lora) prepends LoRA parameter efficient tuning paradigm into LLM. In the training stage, it only trains the linear projector and LoRA parameters, with frozen vision encoder and LLM; MoELoRA (coin) is based on the LoRA, and the number of experts for each MoE layer is set to 2; LWF (lwf) calculates the results of the new dataset samples on both the old and new models. After that, it calculates the distillation loss and adds it to the loss function as a regularization penalty term. EWC (ewc) considers the change of the training parameters and proposes the specific parameters changing loss as a regularization penalty. PGP (pgp) introduces a gradient projection method for efficient parameters, and changes the gradient direction orthogonal to the previous feature subspace. E-Proj (mcit) is the representative SOTA dynamic modal method, which expands the visual projection layer with the task increasing. MT (mt) is the representative SOTA regularization method, which compensates for changes in trainable parameters.
@@
It is worth noting that our judgment of whether the prediction results are correct or not is strictly based on the direct comparison between outputs of MLLMs and ground truths, which is defined as Truth Alignment in (coin). Therefore, our judgment criteria would be more stringent.
@@
New Accuracy (New.ACC) is employed to average the test accuracy of new datasets, which means the plasticity performance.
@@
$\text{Average Accuracy}=\frac{1}{T}\sum_{i=1}^{T}A_{T,i},$ (49)
$\text{Forgetting}=\frac{1}{T-1}\sum_{i=1}^{T-1}{\text{max}(A_{j,i})_{j\in[i,T-1]}-A_{T,i}},$ (50)
$\text{New Accuracy}=\frac{1}{T}\sum_{i=1}^{T}A_{i,i},$ (51)
@@
FT-init 35.70.2 18.50.7 -4.60.2
L2 35.60.1 17.50.5 -3.81.2
EWC 34.50.6 16.80.4 -6.81.5
AGEM (10) 33.20.4 19.10.1 -7.31.0
AGEM (50) 34.90.9 18.11.0 -6.00.9
Replay (10) 38.40.7 23.70.0 -1.30.5
@@
Dataset. We use the InstrDialog Stream dataset from (citb), which consists of 4 tasks from dialogue state tracking, 11 tasks from dialogue generation, and 4 tasks from intent identification, resulting in a total of 19 dialogue tasks.
Model. We use the instruction-tuned T5 model from (citb), which has learned to understand some instructions and can act as a good starting point to conduct subsequent learning.
Comparison Methods. (1). L2 and EWC regularization methods, both of which leverage a Fisher information matrix to mitigate forgetting by regularizing the loss function, thereby penalizing changes to the crucial parameters of previous tasks. (2). Replay method, which stores random instances from each task in a memory buffer, then trains the model jointly on both the new task data and the stored instances. (3). AGEM method, which adds constraints to prevent parameter update from increasing the loss of each previous task. The loss for these tasks is computed based on the instances saved in memory. (4). We adopt AdapterCL, which freezes the pre-trained model and trains an independent Adapter for each task. (5). For ours, we freeze the pre-trained model and train a set of Adapters. Specifically, we utilize the instruction grouping strategy to decide whether to retrain or expand the Adapter according to the current task instructions and adopt dynamical EMA update when retraining the Adapter. Notice here, besides AdapterCL and ours, which freeze the LLM and only train the Adapter, others adopt the full fine-tuning (FFT) mode.
Metrics. Following (citb), we use the following metrics to measure the continual tuning performance. Let $a_{j,i}$ be the ROUGE-L score of the model on the test set of task $t_{i}$ after training on task $t_{j}$, we define the following:
@@
$\textbf{\text{AR}}=\frac{1}{T}\sum_{i=1}^{T}a_{T,i}.$ (52)
@@
$\textbf{\text{FWT}}=\frac{1}{T-1}\sum_{i=2}^{T}a_{i-1,i}.$ (53)
@@
$\textbf{\text{BWT}}=\frac{1}{T-1}\sum_{i=1}^{T-1}\left(a_{T,i}-a_{i,i}\right).$ (54)
@@
In order to validate the robustness of our method, we adopt the following three types of training instructions. For detailed instruction templates please kindly refer to Table 9.
1). Original instruction type: Each task owns only one instruction, and several tasks share the same instructions.
2). Diverse instruction type: Each task owns only one instruction, and different tasks are tailored to distinct instructions.
@@
Appendix P Deduction of high-dimensional $\Theta_{t}$
For high-dimensional $\Theta_{t}$, we have the following Taylor expansion results:
$\mathcal{L}(\Theta)=\mathcal{L}(\Theta_{t})+\left(\frac{\partial\mathcal{L}}{\partial\Theta}\right)^{T}(\Theta-\Theta_{t})+(\Theta-\Theta_{t})^{T}\frac{H}{2}(\Theta-\Theta_{t})+O(\Theta-\Theta_{t})^{3}.$ (55)
$\mathcal{L}(\Theta_{t}^{*})-\mathcal{L}(\Theta_{t})=\left(\frac{\partial\mathcal{L}}{\partial\Theta}\right)^{T}\bigg|_{\Theta=\Theta_{t}^{*}}(\Theta_{t}^{*}-\Theta_{t})+(\Theta_{t}^{*}-\Theta_{t})^{T}\frac{H}{2}(\Theta_{t}^{*}-\Theta_{t}).$ (56)
Notice that we have omitted the high-order infinitesimal term of $O(\Theta-\Theta_{t})^{3}$. In this way, we can obtain the following minimum optimization goal:
$F=\left(\frac{\partial\mathcal{L}}{\partial\Theta}\right)^{T}\bigg|_{\Theta=\Theta_{t}^{*}}(\Theta_{t}^{*}-\Theta_{t})+(\Theta_{t}^{*}-\Theta_{t})^{T}\frac{H}{2}(\Theta_{t}^{*}-\Theta_{t})+e^{T}(\Theta_{t}^{*}-\Theta_{t-1}^{*})+\lambda e^{T}(\Delta\Theta+\Theta_{t-1}^{*}-\Theta_{t}^{*}).$ (57)
The following deduction is similar to the process of $\theta$, finally, we have:
$0=\frac{1}{(\beta_{t}-1)}\left(\frac{\partial\mathcal{L}}{\partial\Theta}\right)^{T}\bigg|_{\Theta=\Theta_{t}^{*}}\Delta\Theta+\frac{\beta_{t}}{(\beta_{t}-1)^{2}}\Delta\Theta^{T}H\Delta\Theta+\frac{e^{T}\Delta\Theta}{\beta_{t}-1}.$ (58)
Considering that $\Delta\Theta=(\Theta_{t-1}^{*}-\Theta_{t})(\beta_{t}-1)$, we can obtain:
$0=\left(\frac{\partial\mathcal{L}}{\partial\Theta}\right)^{T}\bigg|_{\Theta=\Theta_{t}^{*}}(\Theta_{t-1}^{*}-\Theta_{t})+\beta_{t}(\Theta_{t-1}^{*}-\Theta_{t})^{T}H(\Theta_{t-1}^{*}-\Theta_{t})+e^{T}(\Theta_{t-1}^{*}-\Theta_{t}).$ (59)
Obviously, $\beta_{t}$ equals to:
$\beta_{t}=[\left(\frac{\partial\mathcal{L}}{\partial\Theta}\right)^{T}\bigg|_{\Theta=\Theta_{t}^{*}}+e^{T}]\cdot[(\Theta_{t}-\Theta_{t-1}^{*})^{T}H]^{-1}.$ (60)
It can be discovered that Eq.(60) has the same form as Eq.(12). Thus, we can draw the conclusion that our method can be expanded to high-dimensional $\Theta_{t}$.
@@
ScienceQA
Answer with the option’s
letter from the given
choices directly
Answer with the option’s
letter from the given
choices directly
Answer with the option’s letter from the given choices directly
Select the correct answer from the given choices and respond with the letter of the chosen option
Determine the correct option from the provided choices and reply with its corresponding letter
Pick the correct answer from the listed options and provide the letter of the selected option
Identify the correct choice from the options below and respond with the letter of the correct option
From the given choices, choose the correct answer and respond with the letter of that choice
Choose the right answer from the options and respond with its letter
Select the correct answer from the provided options and reply with the letter associated with it
From the given choices, select the correct answer and reply with the letter of the chosen option
Identify the correct option from the choices provided and respond with the letter of the correct option
From the given choices, pick the correct answer and respond by indicating the letter of the correct option
TextVQA
Answer the question
using a single
word or phrase
Capture the essence of your
response in a single word
or a concise phrase
Answer the question with just one word or a brief phrase
Use one word or a concise phrase to respond to the question
Answer using only one word or a short, descriptive phrase
Provide your answer in the form of a single word or a brief phrase
Use a single word or a short phrase to respond to the question
Summarize your response in one word or a concise phrase
Respond to the question using a single word or a brief phrase
Provide your answer in one word or a short, descriptive phrase
Answer the question with a single word or a brief, descriptive phrase
Capture the essence of your response in one word or a short phrase
Capture the essence of your response in a single word or a concise phrase
ImageNet
Answer the object in the
image using a single
word or phrase
Express the object in
the image in
a single word or a
short, descriptive phrase
Summarize the object in the image in a single word or a brief phrase
Provide the object in the image using a single word or a brief phrase
Give the object in the image in the form of a single word or a concise phrase
Express the object in the image with one word or a short, descriptive phrase
Identify the type of content in the image using one word or a concise phrase
Respond to the object in the image with a single word or a short, descriptive phrase
Describe the content of the image using one word or a concise phrase
Express the object in the image in a single word or a short, descriptive phrase
Use a single word or a short phrase to categorize the image content
Classify the image content using only one word or a brief phrase
Use one word or a short phrase to classify the content of the image
GQA
Answer the question
using a single
word or phrase
Respond to the question
briefly, using only one
word or a phrase
Respond to the question with a single word or a short phrase
Respond to the question using only one word or a concise phrase
Answer the question with a single word or a brief phrase
Respond with one word or a short phrase
Provide your answer in the form of a single word or a concise phrase
Respond to the question with just one word or a brief phrase
Answer the question using a single word or a concise phrase
Provide your response using only one word or a short phrase
Respond to the question with a single word or a brief phrase
Respond to the question using just one word or a concise phrase
Answer the question with one word or a short phrase
VizWiz
Answer the question
using a single
word or phrase
Provide a succinct
response with a single
word or phrase
Answer the question using only one word or a concise phrase
Respond to the question using only one word or a concise phrase
Respond to the question with a single word or a brief phrase
Provide your answer using just one word or a short phrase
Respond with one word or a concise phrase
Answer the question with just one word or a brief phrase
Use a single word or a short phrase to answer the question
Provide your answer in the form of one word or a brief phrase
Reply to the question using one word or a concise phrase
Answer with a single word or a short phrase
Use one word or a brief phrase to answer the question
Grounding
Please provide the bounding
box coordinate of the region
this sentence describes
Please provide the bounding
box coordinate of the region
this sentence describes
Identify and provide the bounding box coordinates that match the description given in this sentence
Extract and provide the bounding box coordinates based on the region described in the sentence
Please provide the bounding box coordinate of the region this sentence describes
Find and provide the bounding box coordinates for the region mentioned in the sentence
Provide the coordinates of the bounding box that correspond to the region described in the sentence
Give the bounding box coordinates as described in the sentence
Determine and provide the bounding box coordinates based on the description in the sentence
Identify and provide the coordinates of the bounding box described in the sentence
Provide the coordinates for the bounding box based on the region described in the sentence
Extract and provide the coordinates for the bounding box described in the sentence
Identify and give the coordinates of the bounding box as described by the sentence
VQAv2
Answer the question
using a single
word or phrase
Answer the question
using a single
word or phrase
Answer the question using a single word or phrase
Answer the question with a single word or a brief phrase
Use one word or a short phrase to respond to the question
Answer the question using just one word or a concise phrase
Provide your answer to the question using only one word or a brief phrase
Respond to the question with a single word or a short phrase Use a single word or phrase to answer the question
Provide an answer using only one word or a brief phrase
Answer the question succinctly with one word or a brief phrase
Answer the question with just one word or a short phrase
Respond to the question using a single word or a concise phrase
OCR-VQA
Answer the question
using a single
word or phrase
Condense your answer for
each question into a single
word or concise phrase
Respond to the question with a single word or a short phrase
Answer the question using a single word or a concise phrase
Provide your response using only one word or a short phrase
Use one word or a brief phrase to answer the question
Reply to the question using one word or a concise phrase
Use a single word or a short phrase to answer the question
Use a single word or phrase to answer the question
Provide an answer using only one word or a brief phrase
Provide your answer to the question using only one word or a brief phrase
Respond to the question using a single word or a concise phrase
Answer the question using a single word or phrase